OpenStudy (dumbsearch2):

Solve the following system using any method. 5x - 2y = -16 3x + 5y = -22 (7, 2) (-4,-2) (4, 6) No solution. The system is inconsistent. Infinite number of solutions. The system is dependent.

OpenStudy (dumbsearch2):

Thanks in advance guys!!! :)

OpenStudy (anonymous):

5x - 2y = -16 3x + 5y = -22 OK for this two ways, well actually 3. First, put in x and y into both equations and see if it works. Do you know elimination?

OpenStudy (anonymous):

Or substitution?

OpenStudy (anonymous):

Well... eliminate either x or y by making the values of them in both equations the same. So for this one, I will eliminate x. The LCM of 5 and 3 is 15. In both equations we want 15x. Multiply the first equation by 3, and the second by 5 to get: 15x - 6y = -48 15x + 25y = -110 Now Subtract one from the other: 15x - 6y - (15x + 25y) = -48 -(-110) -31y = 62 y = -2 Now substitute y = -2 into any of the first equations. I'll choose eq 1 3x + 5(-2) = -22 3x - 10 = -22 3x = -12 x = -4 And there you have it

OpenStudy (anonymous):

@dumbsearch2

OpenStudy (dumbsearch2):

Thank you so much!!! :)

OpenStudy (dumbsearch2):

Your AMAZING! :)

OpenStudy (anonymous):

Your welcome. Do you know the other method? Substitution?

OpenStudy (anonymous):

Or you could use matrices...

OpenStudy (anonymous):

You could indeed

OpenStudy (dumbsearch2):

thank you! :)

OpenStudy (dumbsearch2):

so much for your help! :)

OpenStudy (dumbsearch2):

matrices? what dat? @Dido525

OpenStudy (anonymous):

Never mind then :P . Wait till linear algebra.